# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v4

@@ 9.5 The index pairing for the reduced group $\mathrm{C}^{*}$-algebra of the braid group $\mathrm{B}_{3}$
The 2-positive elements are precisely the elements $\sigma_{2}^{k}$ with $k\geqslant 0$. Examples of 1-positive element are given by $\sigma_{2}\sigma_{1}^{k}\sigma_{2}^{-2}$ or $\sigma_{1}^{k}$ with $k\geqslant 0$.
The K-theory groups of the braid group $\mathrm{B}_{3}$ are known. Actually, by [ABARW22, Proposition 6.4 p. 1288], we have isomorphisms
$\mathrm{K}_{0}(\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{3}))=\mathrm{K}_{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{3}))=\mathbb{Z}.$
Moreover, the group $\mathrm{K}_{0}(\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{3}))$ is generated by the unit of $\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{3})$ and the group $\mathrm{K}_{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{3}))$ is generated by $\lambda_{\sigma_{1}}]=[\lambda_{\sigma_{2}}]$. In the next result, we compute the index pairing of (9.6).
@@
Suppose that $1<p<\infty$. For any integer $k\in\mathbb{Z}$, we have
(9.11) $\big\langle[\lambda_{\sigma_{1}^{k}}],(\mathrm{L}^{p}(\mathrm{VN}(\mathrm{B}_{3})),F)\big\rangle_{\mathrm{K}_{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{3})),\mathrm{K}^{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{3})),\mathscr{L}^{p}_{\mathrm{nc}})}=-k.$
Proof : If $i=1$ or $i=2$, we have for any $s\in\mathrm{B}_{3}$
$T_{\lambda_{\sigma_{i}^{k}}}(\lambda_{s})=P_{+}(\lambda_{\sigma_{i}^{k}}\lambda_{s})=P_{+}(\lambda_{\sigma_{i}^{k}s}).$
Suppose that $k\geqslant 0$. If $s$ is 1-positive then $\lambda_{\sigma_{1}^{k}s}$ is also 1-positive, so $T_{\lambda_{\sigma_{1}^{k}}}(\lambda_{s})=\lambda_{\sigma_{1}^{k}s}$. If $s$ is 2-positive then $\lambda_{\sigma_{2}^{k}s}$ is also 2-positive. Hence $T_{\lambda_{\sigma_{2}^{k}}}(\lambda_{s})=\lambda_{\sigma_{2}^{k}s}$. If $s=e$ then $T_{\lambda_{\sigma_{2}^{k}}}(e)=\lambda_{\sigma_{2}^{k}}$. By linearity, it is easy to check that $\operatorname{Ker}T_{\lambda_{\sigma_{1}^{k}}}=\{0\}$. Consequently, we have
@@
$=-\dim\mathrm{L}^{p}(\mathrm{VN}(\mathrm{B}_{3}))/\operatorname{Ran}T_{\lambda_{\sigma_{1}^{k}}}=-k.$
We conclude with (9.6).
Now, suppose that $k<0$. If $s=\sigma_{2}^{r}$ is $2$-positive then $\sigma_{2}^{k}s=\sigma_{2}^{k}\sigma_{2}^{r}=\sigma_{2}^{r+k}$ is 2-positive if and only if $r+k\geqslant 0$. It is obvious that the left multiplication by $\sigma_{2}^{k}$ transforms any 1-positive element into a 1-positive element. So, it is easy to check that $\dim\operatorname{Ker}T_{\lambda_{\sigma_{2}^{k}}}=k$. We conclude that
@@
In [LOS21, Example 5.10], the $\mathrm{K}$-theory of the reduced group $\mathrm{C}^{*}$-algebra of the group $\mathrm{B}_{4}$ is computed to be
$\mathrm{K}_{0}(\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{4}))\simeq\mathbb{Z}\oplus(\mathbb{Z}/2\mathbb{Z}),\qquad\qquad\mathrm{K}_{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{4}))\simeq\mathbb{Z}.$
The abelian group $\mathrm{K}_{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathrm{B}_{4}))$ is generated by the element $\lambda_{\sigma_{1}}]=[\lambda_{\sigma_{2}}]=[\lambda_{\sigma_{3}}]$. It is not difficult to compute the index pairing.
9.6 Bi-orderings of the Thompson group $\mathrm{F}$
The Thompson group $F$ is the group of homeomorphisms $g:[0,1]\to[0,1]$ for which:
  • the endpoints satisfy $g(0)=0$, $g(1)=1$,
  • it is piecewise linear with finitely many break points on the diadic numbers $\mathbb{Z}[\tfrac{1}{2}]\cap(0,1)$,
  • all slopes of $g$ are in the set $2^{\mathbb{Z}}:=\{2^{n}\mid n\in\mathbb{Z}\}$.
The Thompson group $\mathrm{F}$ consists of orientation-preserving piecewise linear homeomorphisms $f$ of the interval $[0,1]$ such that
  1. 1.
    there are only finitely many points of non-differentiability of $f$,
  2. 2.
    each point of non-differentiability of $f$ occurs at a dyadic rational number $\frac{a}{2^{n}}$ where $a,n\in\mathbb{N}$,
  3. 3.
    the derivatives on intervals of differentiability are powers of 2.
See Figure 1 for an example of element of this group.
Figure 1: an element of Thompson group $\mathrm{F}$
We refer to [CFP96] for an introduction to this group. $F$ is a countable group, and it is generated by the elements $A,B$ whose graphs are shown in Fig. 2. Moreover, it has a finite presentation in terms of $A,B$, namely
$F=\langle A,B\mid[AB^{-1},A^{-1}BA]=[AB^{-1},A^{-2}BA^{2}]=e\rangle,$
where the group commutator $g,h]=ghg^{-1}h^{-1}$ is defined as usual.
Figure 2: the generators $A,B$ of $F$.
A description of all bi-orderings on Thompson group is given in the paper [NaR10]. We describe four of them. Let $f\in\mathrm{F}$ with $f\not=\mathrm{Id}$. We consider the corresponding lateral derivatives $f^{\prime}_{-}$ and $f^{\prime}_{+}$. We denote by $x^{-}_{f}$ the leftmost point $x$ for which $f^{\prime}_{+}(x)\neq 1$ and we denote by $x^{+}_{f}$ the rightmost point $x$ for which $f^{\prime}_{-}(x)\neq 1$. In this context, we can define
  1. 1.
    the bi-ordering $\preceq_{x^{-}}^{+}$ for which $\mathrm{Id}\prec f$ if and only if $f^{\prime}_{+}(x^{-}_{f})>1$,
  2. 2.
    the bi-ordering $\preceq_{x^{-}}^{-}$ for which $\mathrm{Id}\prec f$ if and only if $f^{\prime}_{+}(x^{-}_{f})<1$,
  3. 3.
    the bi-ordering $\preceq_{x^{+}}^{+}$ for which $\mathrm{Id}\prec f$ if and only if $f^{\prime}_{-}(x^{+}_{f})<1$,
  4. 4.
    the bi-ordering $\preceq_{x^{+}}^{-}$ for which $\mathrm{Id}\prec f$ if and only if $f^{\prime}_{-}(x^{+}_{f})>1$.
Our results pave the way for a better understanding of the K-theory and K-homology groups of Thompson group, whose properties remain largely unknown.
10 Banach Fredholm module associated to the free Hilbert transform
Let $G=\mathbb{F}_{\infty}$ be the free group with a countable sequence of generators $g_{1},g_{2},\ldots$. For any integer $n\geqslant 1$, let $L_{n}^{+}\colon\mathrm{L}^{2}(\mathrm{VN}(\mathbb{F}_{\infty}))\to\mathrm{L}^{2}(\mathrm{VN}(\mathbb{F}_{\infty}))$ and $L_{n}^{-}\colon\mathrm{L}^{2}(\mathrm{VN}(\mathbb{F}_{\infty}))\to\mathrm{L}^{2}(\mathrm{VN}(\mathbb{F}_{\infty}))$ be the orthogonal projections such that
[…]
Let us further consider signs $\varepsilon_{n}^{+},\varepsilon_{n}^{-}\in\{-1,1\}$ for any integer $n\geqslant 1$. Following [MeR17], we define the free Hilbert transform associated with $\varepsilon=(\varepsilon_{n}^{\pm})$ as
Clearly, since the ranges of the various projections $L_{n}^{\pm}$ are mutually orthogonal, the linear map $\mathcal{H}_{\varepsilon}$ is bounded on the Hilbert space $\mathrm{L}^{2}(\mathrm{VN}(\mathbb{F}_{\infty}))$. The far-reaching generalization in [MeR17, Section 4] is that the map $\mathcal{H}_{\varepsilon}$ induces a completely bounded map on the Banach space $\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))$ for any $1<p<\infty$.
In the sequel, we consider the contractive homomorphism $\pi\colon\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{\infty})\to\mathrm{B}(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))$, $a\mapsto M_{a}$, where $M_{a}\colon\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))\to\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))$, $y\mapsto ay$ is the multiplication operator by $a$.
Proposition 10.1
Suppose that $1<p<\infty$. Consider a family $\varepsilon=(\varepsilon_{n}^{\pm})_{n\geqslant 1}$ of signs. Then $(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty})),\mathcal{H}_{\varepsilon})$ is a Banach Fredholm module over the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{\infty})$.
Proof : Let $s,t\in\mathbb{F}_{\infty}$. Assume that $s$ begins with $g_{k}^{\eta_{s}}$ and that $ts$ begins with $g_{l}^{\mu_{ts}}$, where $\eta_{s},\mu_{ts}\in\{-1,1\}$ (identified with $\{-,+\}$). We have
$\left[\mathcal{H}_{\varepsilon},M_{\lambda_{t}}\right](\lambda_{s})=\mathcal{H}_{\varepsilon}M_{\lambda_{t}}(\lambda_{s})-M_{\lambda_{t}}\mathcal{H}_{\varepsilon}(\lambda_{s})$
[…]
$=\varepsilon_{l}^{\mu_{ts}}\lambda_{ts}-M_{\lambda_{t}}(\varepsilon_{k}^{\eta_{s}}\lambda_{s})=\varepsilon_{l}^{\mu_{ts}}\lambda_{ts}-\varepsilon_{k}^{\eta_{s}}\lambda_{ts}.$
The other cases where $s=e$ or $ts=e$ are left to the reader. We deduce that the rank of the commutator $\left[\mathcal{H}_{\varepsilon},M_{\lambda_{t}}\right]$ is zero or one. If $a=\sum_{s\in\mathbb{F}_{\infty}}a_{s}\lambda_{s}$ is a finite sum, where each $a_{s}$ is a complex number, belonging to the group algebra $\mathbb{C}[\mathbb{F}_{\infty}]$ then we see by linearity that the commutator $\left[\mathcal{H},M_{a}\right]$ is a finite-rank operator. We infer that for any $a\in\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{\infty})$ the commutator $\left[\mathcal{H},M_{a}\right]$ is compact, as a limit of finite-rank operators, since the homomorphism $\pi\colon\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{\infty})\to\mathrm{B}(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty})))$, $a\mapsto M_{a}$ is continuous.
Note that $\mathcal{H}_{\varepsilon}(\lambda_{e})=0$. Consequently, with (10.1) it is easy to check that $\mathcal{H}_{\varepsilon}^{2}\sim\mathrm{Id}_{\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty})}$.
Remark 10.2
By restriction, we can obtain a Banach Fredholm module over each reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n})$ for any integer $n\geqslant 1$. By [PiV82, Corollary 3.2 p. 152], we have $\mathrm{K}_{0}(\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n}))=\mathbb{Z}$ and the generator of $\mathrm{K}_{0}(\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n}))$ is $[1]$. The same reference gives an isomorphism $\mathrm{K}_{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n}))=\mathbb{Z}^{n}$ and the generators of the group $\mathrm{K}_{1}(\mathrm{C}^{*}_{\lambda}(\mathbb{F}_{n}))$ are $\lambda_{g_{1}}],\ldots,[\lambda_{g_{n}}]$. We invite the reader to determine the index pairing.

### v5

@@ 9.5 Banach Fredholm module associated to the free Hilbert transform
The ranges of the various projections $L_{n}^{\pm}$ are mutually orthogonal. Let us further consider signs $\varepsilon_{n}^{+},\varepsilon_{n}^{-}\in\{-1,1\}$ for any integer $n\geqslant 1$. Following [MeR17], we define the free Hilbert transform associated with the family $\varepsilon=(\varepsilon_{n}^{\pm})_{n\geqslant 1}$ of signs by
For any $s\neq e$, observe that
(9.9) $\mathcal{H}_{\varepsilon}(\lambda_{s})=\pm\lambda_{s}\quad\text{and}\quad\mathcal{H}_{\varepsilon}(1)=0.$
By [MeR17], the map $\mathcal{H}_{\varepsilon}$ extends to a completely bounded operator on the space $\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))$ for any $1<p<\infty$.
For any $a\in\mathrm{VN}(\mathbb{F}_{\infty})$, we denote by $R_{a}\colon\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))\to\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))$, $x\mapsto xa$ the right multiplication operator. Then by Hölder’s inequality we have $\left\|R_{a}\right\|_{\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))\to\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))}\leqslant\left\|a\right\|_{\mathrm{VN}(\mathbb{F}_{\infty})}$. So the linear map $\mathrm{VN}(\mathbb{F}_{\infty})\to\mathrm{B}(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty})))$, $a\mapsto R_{a}$ is a contractive homomorphism and restricts to a contractive homomorphism $\pi$ of $\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{\infty})$ into the space $\mathrm{B}(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty})))$.
@@
Suppose that $1<p<\infty$. Consider a family $\varepsilon=(\varepsilon_{n}^{\pm})_{n\geqslant 1}$ of signs. Then $(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty})),\pi,\mathcal{H}_{\varepsilon})$ is a Banach Fredholm module over the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{\infty})$.
@@
Fix an element $t\in\mathbb{F}_{\infty}$ and write its reduced form as
$t=a_{1}a_{2}\cdots a_{m},$
where each $a_{j}$ is one of the letters $g_{n}$ or $g_{n}^{-1}$. For any integer $r\in\{0,\ldots,m\}$, set
$t_{[1,r]}\overset{\mathrm{def}}{=}a_{1}a_{2}\cdots a_{r},$
with the convention $t_{[1,0]}\overset{\mathrm{def}}{=}e$. Define the finite subset
$E_{t}\overset{\mathrm{def}}{=}\big\{(t_{[1,r]})^{-1}:0\leqslant r\leqslant m\big\}$
of the group $\mathbb{F}_{\infty}$. Let $s\in\mathbb{F}_{\infty}$ and consider the reduced product $st$. Cancellation between $s$ and $t$ can only occur at the junction between the last letters of $s$ and the first letters of $t$. Hence the first letter of the reduced word of $st$ coincides with the first letter of $s$, except possibly when $s$ is entirely cancelled by an initial segment of $t$. Equivalently, the first letter of $st$ differs from that of $s$ only when $s\in E_{t}$. In particular, if $s\notin E_{t}$ we have
(9.10) $\mathcal{H}_{\varepsilon}(\lambda_{s})=\sigma(s)\lambda_{s},\qquad\mathcal{H}_{\varepsilon}(\lambda_{st})=\sigma(s)\lambda_{st}$
for the same sign $\sigma(s)\in\{-1,1\}$ determined by the first letter of $s$. Note that $e\in E_{t}$, hence the case $s=e$ is included in the exceptional set and does not require a sign $\sigma(s)$. We deduce that
$\left[\mathcal{H}_{\varepsilon},R_{\lambda_{t}}\right](\lambda_{s})=\mathcal{H}_{\varepsilon}(R_{\lambda_{t}}\lambda_{s})-R_{\lambda_{t}}\mathcal{H}_{\varepsilon}(\lambda_{s})=\mathcal{H}_{\varepsilon}(\lambda_{st})-\mathcal{H}_{\varepsilon}(\lambda_{s})\lambda_{t}$
@@
Therefore the commutator vanishes on the linear span of the subset $\{\lambda_{s}:s\notin E_{t}\}$ and its range is contained in the finite-dimensional space
$\operatorname{span}\{\lambda_{st}:s\in E_{t}\}.$
Hence the commutator $\mathcal{H}_{\varepsilon},R_{\lambda_{t}}]$ has finite rank on the space $\mathbb{C}[\mathbb{F}_{\infty}]$ and extends by continuity to a finite-rank operator on the Banach space $\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))$.
Let $a\in\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{\infty})$ and choose a sequence $(a_{k})_{k\geqslant 1}$ in the group algebra $\mathbb{C}[\mathbb{F}_{\infty}]$ such that $a_{k}\to a$ in $\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{\infty})$. We can write each $a_{k}$ as a finite sum $a_{k}=\sum_{t\in\mathbb{F}_{\infty}}a_{k,t}\lambda_{t}$ where $a_{k,t}\in\mathbb{C}$. Consequently, we have
$\mathcal{H}_{\varepsilon},R_{a_{k}}]=\sum_{t\in\mathbb{F}_{\infty}}a_{k,t}[\mathcal{H}_{\varepsilon},R_{\lambda_{t}}]$
is a finite sum of finite-rank operators, hence finite-rank and therefore compact on the Banach space $\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{\infty}))$. Since the map $a\mapsto R_{a}$ is continuous on the space $\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{\infty})$, we have $R_{a_{k}}\to R_{a}$. So, we obtain
$\left\|[\mathcal{H}_{\varepsilon},R_{a_{k}}]-[\mathcal{H}_{\varepsilon},R_{a}]\right\|\leqslant 2\left\|\mathcal{H}_{\varepsilon}\right\|\left\|R_{a_{k}}-R_{a}\right\|\xrightarrow[k\to\infty]{}0.$
We infer that the commutator $\mathcal{H}_{\varepsilon},R_{a}]$ is a norm-limit of compact operators, hence compact.
By restriction, we can obtain a Banach Fredholm module over each reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n})$ for any integer $n\geqslant 1$. Recall that by [PiV82, Corollary 3.2 p. 152], we have $\mathrm{K}_{0}(\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n}))=\mathbb{Z}$ and the generator of $\mathrm{K}_{0}(\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n}))$ is $[1]$. The same reference gives an isomorphism $\mathrm{K}_{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n}))=\mathbb{Z}^{n}$ and the generators of the group $\mathrm{K}_{1}(\mathrm{C}^{*}_{\lambda}(\mathbb{F}_{n}))$ are $\lambda_{g_{1}}],\ldots,[\lambda_{g_{n}}]$.
Proposition 9.7
Suppose that $1<p<\infty$. Let $\varepsilon=(\varepsilon_{j}^{\pm})_{1\leq j\leq n}$ be a family of signs. Consider the odd Banach Fredholm module $(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{n})),\pi,\mathcal{H}_{\varepsilon})$ over the algebra $\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n})$. Then, for any integer $j\in\{1,\ldots,n\}$, the index pairing of the element $\lambda_{g_{j}}]\in\mathrm{K}_{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n}))$ with the class $(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{n})),\pi,\mathcal{H}_{\varepsilon})]\in\mathrm{K}^{1}(\mathrm{C}^{*}_{\mathrm{red}}(\mathbb{F}_{n}),\mathscr{B})$ is given by
(9.11) $\bigl\langle[\lambda_{g_{j}}],[(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{n})),\pi,\mathcal{H}_{\varepsilon})]\bigr\rangle=\delta_{\varepsilon_{j}^{-}=1}-\delta_{\varepsilon_{j}^{+}=1}.$
In particular, if $\varepsilon_{j}^{+}=1$ and $\varepsilon_{j}^{-}=-1$, then $\langle[\lambda_{g_{j}}],[(\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{n})),\pi,\mathcal{H}_{\varepsilon})]\rangle=-1$.
Proof : We use Theorem 4.1 with $u=\lambda_{g_{j}}$ and $n=1$. Set $F\overset{\mathrm{def}}{=}\mathcal{H}_{\varepsilon}$ and $u_{1}\overset{\mathrm{def}}{=}\pi(\lambda_{g_{j}})=R_{\lambda_{g_{j}}}$. The Fredholm operator of Theorem 4.1 is
$T=Pu_{1}P-(\mathrm{Id}-P),\qquad P=\frac{\mathrm{Id}+F}{2}.$
Since $F(1)=0$, we have $P(1)=\frac{1}{2}1$ and $P$ is not a projection. Now, we replace $P$ by an idempotent $P_{0}$ differing from $P$ by a finite-rank operator. Define the bounded operator $P_{0}\colon\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{n}))\to\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{n}))$ by $P_{0}\overset{\mathrm{def}}{=}P-\frac{1}{2}Q$. We have
$P_{0}(1)=0,\qquad P_{0}(\lambda_{s})=\begin{cases}\lambda_{s}&\text{if }s\neq e\text{ and }F(\lambda_{s})=+\lambda_{s},\\ 0&\text{if }s\neq e\text{ and }F(\lambda_{s})=-\lambda_{s}.\end{cases}$
Then $P_{0}$ is an idempotent and $P-P_{0}$ has rank one and in particular is compact. Set
$T_{0}\overset{\mathrm{def}}{=}P_{0}u_{1}P_{0}-(\mathrm{Id}-P_{0}).$
Since $u_{1}$ is bounded, we have
$T-T_{0}=(P-P_{0})u_{1}P+P_{0}u_{1}(P-P_{0})+(P-P_{0}),$
which is a sum of products involving the compact operator $P_{0}-P$ and bounded operators. Hence $T-T_{0}$ is compact and therefore
[…]
Let $X_{+}\overset{\mathrm{def}}{=}\operatorname{Ran}P_{0}$ and $X_{-}\overset{\mathrm{def}}{=}\operatorname{Ran}(\mathrm{Id}-P_{0})$. Then $\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{n}))=X_{+}\oplus X_{-}$ topologically and $P_{0}$ is the projection onto $X_{+}$ along $X_{-}$. By construction, $X_{+}$ is the closed linear span of
[…]
Consider the operator
$S\overset{\mathrm{def}}{=}P_{0}u_{1}P_{0}\colon X_{+}\to X_{+}.$
In the decomposition $\mathrm{L}^{p}(\mathrm{VN}(\mathbb{F}_{n}))=X_{+}\oplus X_{-}$, the operator $T_{0}$ acts as $S$ on $X_{+}$ and as $\mathrm{Id}$ on $X_{-}$. Hence $T_{0}$ is Fredholm if and only if $S$ is Fredholm and
(9.13) $\operatorname{Index}T_{0}=\operatorname{Index}S.$
Let $x\in X_{+}$. Then $Sx=0$ means $P_{0}(x\lambda_{g_{j}})=0$. Since $P_{0}$ is diagonal on the Fourier basis, it suffices to test basis vectors. Let $s\neq e$ with $\lambda_{s}\in X_{+}$. Then $S(\lambda_{s})=0$ if and only if $\lambda_{sg_{j}}\in X_{-}$, namely if the reduced word of $sg_{j}$ starts with a letter $g_{k}^{\pm 1}$ with $\varepsilon_{k}^{\pm}=-1$, or if $sg_{j}=e$.
Now $sg_{j}=e$ if and only if $s=g_{j}^{-1}$. In that case $S(\lambda_{g_{j}^{-1}})=P_{0}(1)=0$. Moreover, $\lambda_{g_{j}^{-1}}\in X_{+}$ if and only if $\varepsilon_{j}^{-}=+1$. If $s\neq g_{j}^{-1}$, then the first letter of $sg_{j}$ equals the first letter of $s$, so $\lambda_{sg_{j}}\in X_{+}$ whenever $\lambda_{s}\in X_{+}$, hence $S(\lambda_{s})v\neq 0$. Therefore
$\operatorname{Ker}S=\begin{cases}\operatorname{span}\lambda_{g_{j}^{-1}}&\text{if }\varepsilon_{j}^{-}=+1,\\ \{0\}&\text{if }\varepsilon_{j}^{-}=-1.\end{cases}$
In particular, we have
(9.14) $\dim\operatorname{Ker}S=\delta_{\varepsilon_{j}^{-}=+1}.$
Now, we claim that
$X_{+}/\operatorname{Ran}S\cong\begin{cases}\mathbb{C}\lambda_{g_{j}}&\text{if }\varepsilon_{j}^{+}=+1,\\ \{0\}&\text{if }\varepsilon_{j}^{+}=-1,\end{cases}$
and hence
(9.15) $\mathrm{codim}\,\operatorname{Ran}S=\delta_{\varepsilon_{j}^{+}=+1}.$
Indeed, fix a basis vector $\lambda_{w}\in X_{+}$. If $w\neq g_{j}$, then $w$ admits a reduced form not equal to the one-letter word $g_{j}$. Let $s=wg_{j}^{-1}$ reduced. Then $s\neq e$ and the reduced word of $s$ starts with the same first letter as $w$. Since $\lambda_{w}\in X_{+}$, this first letter has positive sign for $\varepsilon$, hence $\lambda_{s}\in X_{+}$. Moreover, we have
$S(\lambda_{s})=P_{0}(\lambda_{sg_{j}})=P_{0}(\lambda_{w})=\lambda_{w},$
so $\lambda_{w}\in\operatorname{Ran}S$. Therefore every basis vector of $X_{+}$ except possibly $\lambda_{g_{j}}$ belongs to $\operatorname{Ran}S$.
Finally, consider $\lambda_{g_{j}}$. If $\varepsilon_{j}^{+}=-1$, then $\lambda_{g_{j}}\notin X_{+}$ and there is nothing to prove. If $\varepsilon_{j}^{+}=+1$, then $\lambda_{g_{j}}\in X_{+}$. Suppose by contradiction that $\lambda_{g_{j}}\in\operatorname{Ran}S$. Then there exists $x\in X_{+}$ with $Sx=\lambda_{g_{j}}$, meaning $P_{0}(x\lambda_{g_{j}})=\lambda_{g_{j}}$. Writing $x=\sum_{s}c_{s}\lambda_{s}$ as a finite sum in $\mathbb{C}[\mathbb{F}_{n}]$ first, we see that the coefficient of $\lambda_{g_{j}}$ in $x\lambda_{g_{j}}$ is exactly the coefficient of $\lambda_{e}$ in $x$. But $x\in X_{+}$ implies $P_{0}(x)=x$ and $P_{0}(1)=0$, hence the coefficient of $\lambda_{e}$ in $x$ is 0. Therefore the coefficient of $\lambda_{g_{j}}$ in $x\lambda_{g_{j}}$ is 0, a contradiction. By density, the same conclusion holds for general $x\in X_{+}$. Thus $\lambda_{g_{j}}\notin\operatorname{Ran}S$ when $\varepsilon_{j}^{+}=+1$, and (9.15) follows.
We conclude that
[…]
This equals the index pairing given by (4.3), and (9.11) follows.
@@
This result and its proof show that it provides a generalization of Gohberg-Krein index theorem, described in (8.20).